OpenStudy (anonymous):

If 3a-c=5b and 3a+3b-c=40, what is the value of b?

OpenStudy (anonymous):

We have: \[ 3a-c=5b \]And: \[ 3a+3b-c=40\implies\\ (3a-c)+3b=40\implies\\ 5b+3b=40\implies\\ 8b=40\implies\\ b=5 \]And we're done.